OpenStudy (anonymous):

the height of a hexagonal prism is 6ft. Its base edges all equal 4ft. Find its volume

OpenStudy (lgbasallote):

hmm the figures are getting harder...

OpenStudy (anonymous):

yeah i never learned this -_- dang online school ...

OpenStudy (lgbasallote):

yeahh they never teach anything..just homeworks and seatworks then quiz

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

The formula for calculating the volume of a hexagonal prism is to take the area of the hexagon, then multiply it by the height of the prism.

OpenStudy (lgbasallote):

first thing you need to know is that the volume of prisms is always \[V = \text{base area} \times \text{height}\]so yeah you need to find the area of the hexagon first

OpenStudy (anonymous):

Everything is in your online school's lessons by the way. Don't ask people for their answers, just go back into the lesson.

OpenStudy (anonymous):

Volume = 1/2 * apathem * number of sides * side length * height number of sides: 6 side lengt: 4 ft height: 6 ft apthem: ? so we are missing the apathem Apathem is the segment from the center perpendicular to a side and bisect that side. so [ draw an apathem. It'll form 2 small right triangle, with the short legs of 2ft. The apathem is the long leg of one of the small triangles. Since the 2 small right triangles are special angles 30-60-90. The apthem is 2sqrt(3) Volume = 1/2 * 2sqrt(3) * 6 * 4 * 6 Volume = 249.41 ft^3 o and do i win the medal :)

OpenStudy (anonymous):

im not asking for there answers im asking for there help ... me i need someone to explain it to me thats how i learn i know the lessons are in the online thing but i dont understand it cause i cant just look at it and figure it out everyone learns different

OpenStudy (anonymous):

Everything was stated for you. Now you should have all the information to solve this on your own.

OpenStudy (anonymous):

yeah i know now i understand it im not saying i dont .. I just need help thats all. Im not as smart as you buddy haha im dumb in math im not afraid to say it lol

OpenStudy (anonymous):

Not trying to be disrespectful

OpenStudy (anonymous):

I'm terrible with anything Science related. Sorry about the mean remark.

OpenStudy (anonymous):

this is it Volume = 1/2 * 2sqrt(3) * 6 * 4 * 6 Volume = 249.41 ft^3

OpenStudy (anonymous):

Its fine @Dubstep97

OpenStudy (anonymous):

@85295james there are no decimals here

OpenStudy (anonymous):

in the answers

OpenStudy (zzr0ck3r):

The hexagon has 6 sides and 6 vertices. Label the vertices A, B, C, D, E, and F. Draw a line from A → D, B → E, and C→ F. The lines intersect at the center of the hexagon. You should see 6 triangles. The area of the hexagon = 6 * area of 1 triangle. then multiply that by 6

OpenStudy (lgbasallote):

@eAnthonyR do you remember that formula i gave you about the area of equilateral triangle?

OpenStudy (anonymous):

s^2 one?

OpenStudy (lgbasallote):

\[A = \frac{s^2 \sqrt 3}{4}\] a hexagon is comprised of 6 equilateral triangles so area of hexagon is \[A = \frac{s^2 \sqrt 3}{4} \times 6 \implies \frac{3s^2 \sqrt 3}{2}\] so volume is area x height so \[V = \frac{3s^2h \sqrt 3}{2}\] does that make sense? lol

OpenStudy (anonymous):

ok thin it =72 times sqrt 12 or it = 144 times sqrt 3

OpenStudy (anonymous):

oh ok i got it now @lgbasallote And thanks @85295james thats what i got xD

OpenStudy (lgbasallote):

lol cool. i was able to derive the formula of area of hexagon without using apothems =))

OpenStudy (anonymous):

yeah i never learned apothems haha i never heard of it xD